Question 4 [7]


Prove that 𝑓(π‘₯) = π‘₯^2 + 2π‘₯ is not injective.

1
Expert's answer
2022-03-09T10:14:39-0500

Since "𝑓(0) = 0^2 + 2\cdot 0=0" and "\ud835\udc53(-2) = (-2)^2 + 2(-2)=4-4=0,"we conclude that "f(0)=f(-2)," and hence "f" is not injective.